OpenStudy (anonymous):

Evaluate -x^4+3x^1-4 for x=2.

OpenStudy (anonymous):

@NrGized

OpenStudy (anonymous):

@nerdygirl2345

OpenStudy (anonymous):

-(2)^4+3(2)^1-4. 16+6-4 22-4 =18

OpenStudy (anonymous):

I think thats correct. o.o haha

OpenStudy (anonymous):

no ughh but i got the same thing omg

OpenStudy (anonymous):

wait is the equation \[-x^4 + 3x^{1 - 4} =?\]

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

is that the equation?

OpenStudy (anonymous):

-x^4+3x^1-4 for x=2.

OpenStudy (anonymous):

is my equation correct?

OpenStudy (anonymous):

that i made?

OpenStudy (anonymous):

no the equation is -x to the fourth power plus 3x to the first power minus 4

OpenStudy (anonymous):

nope it doesn't make sense

OpenStudy (anonymous):

haha ig

OpenStudy (anonymous):

the answer im getting is not correct then that means that the question is incorrectly written

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

take a picture of the question and then post it on here

OpenStudy (tkhunny):

-x^4+3x^1-4 for x=2.